OpenStudy (anonymous):

Find the Surface Area of a cone with a diameter of 10 cm and a height of 12 cm. Use pi = 3.14. Show your work for credit. Find the Surface Area of a cone with a diameter of 10 cm and a height of 12 cm. Use pi = 3.14. Show your work for credit. @Mathematics

OpenStudy (anonymous):

Would you give me credit too?? :D

hero (hero):

I wouldn't

OpenStudy (anonymous):

You miser!

hero (hero):

\[SA = \pi rs+\pi r^2\]

OpenStudy (anonymous):

lozl toooo cute @ foolformath

hero (hero):

You call that cute?

hero (hero):

Far From it. It's ugly

OpenStudy (anonymous):

Thanks ... you have got a nice name too.. "KISS" is my favorite math technique ;) KISS -> keep it simple silly ;)

OpenStudy (anonymous):

i thought it was cute :/

hero (hero):

I thought it was ugly. Same difference

OpenStudy (anonymous):

Hero is a cartoon you may ignore his opinion .. he is just here to entertain us with his/her sarcasm ..

OpenStudy (anonymous):

:'( jeez hero tht wasnt nice

hero (hero):

Other guy isn't being so nice either. You don't understand

OpenStudy (anonymous):

Lolz and hero have all the intelligence of this planet

OpenStudy (anonymous):

o.O soooo im jus gunna say u both are kool :D haha

OpenStudy (anonymous):

Lolz I am always cool .. I don't about him/her :P Coming back to your question .. you mean total surface area or curved surface area ? :)

OpenStudy (anonymous):

the total surface area :DD haha please

hero (hero):

You're the Master of The Universe and you don't care about me? I'm not surprised.

OpenStudy (anonymous):

huh???

OpenStudy (anonymous):

anyone there :'( ?

OpenStudy (anonymous):

Yes yes .. :) You could do it using that formula hero posted earlier in this thread :)

OpenStudy (anonymous):

OOOOOOOOOOOhhhhhh hahahahaha lozl i missed that completely

OpenStudy (anonymous):

Or do you need more explanation? HINT:"r" is the radius and "s" is the slant height.

OpenStudy (anonymous):

what is slant height? :(

hero (hero):

r^2 + h^2 = s^2

hero (hero):

s = slant height

OpenStudy (anonymous):

pythag theorem ?

OpenStudy (anonymous):

h is the perpendicular distance between the vertex and the center of the circular base of the cone.

hero (hero):

yes, pythagorean theorem

OpenStudy (anonymous):

5^2 + 12^2 = s^2 25 + 144 = s^2 169=s^2 s=13 ?

OpenStudy (anonymous):

Yes you are awesome ^^

hero (hero):

Looks right

OpenStudy (anonymous):

yaaayy :DDD haha so 13 is my answer?

hero (hero):

I'm so glad I didn't have to do anything except argue with FOOL

OpenStudy (anonymous):

It's not just looks right .. it is right !

OpenStudy (anonymous):

yaaay thank u foolformath haha oh wait i have to put tht into the other formula right?

OpenStudy (anonymous):

Well you are not capable of anything else,but I know how to ignore cartoons ;)

hero (hero):

I knew it was right

OpenStudy (anonymous):

yes ... substitute in the total surface area formula and you will have your final answer.

OpenStudy (anonymous):

ok :) i will do it now and post will u chek my work ?

OpenStudy (anonymous):

sure :) also tell me what does kiss69fist23 means ? :P

OpenStudy (anonymous):

SA=πrs+πr2 sa = 3.14(5)13 + 3.14(5)^2 sa = 204.1 + 28.14 sa = 232.24 ??

OpenStudy (anonymous):

oh haha its just a thing i do my email is kiss with a fist .Like kissing with a fist (punch in the mouth) 23 is my fave number nd im perverted so i had to add a 69 hahaha ;) its just a lil piece of me :)

OpenStudy (anonymous):

\[ 3.14 \times 5 \times 13 + 3.14 \times 5^2 = 282.6 \]

OpenStudy (anonymous):

Lolz your bf have no teeth left in his mouth I guess :P

OpenStudy (anonymous):

hahahahah !!